OpenStudy (anonymous):

A boy and girl are balanced on a "massless" seesaw. The boy has a mass of 75 kg and the girl a mass of 50kg. The seesaw is 6 meters in lenght with the pivot point between them...What are the distances each child must sit relative to the pivot point?

OpenStudy (amistre64):

what are your thoughts about it?

OpenStudy (amistre64):

spose the boy is at 1 meter from the pivot, how far does the girl have to be to balance out?

OpenStudy (anonymous):

1 meter from the pivot

OpenStudy (amistre64):

no, a 50 girl does not balance out a 75 boy when they are equal distances from the pivot.

OpenStudy (amistre64):

the force applied at a distance, is equal to Fd 75*1 = 50*d when d=1, we have 75 = 50 which is just not correct

OpenStudy (amistre64):

what value of d, makes 75 = 50d ?

OpenStudy (anonymous):

25

OpenStudy (amistre64):

50*25 is not 75, 50 + 25 is.... but that is not the operation we are doing 50*75/50 = 75 so for every 1 meter the boy sits, the girl has to sit 75/50 meters to balance out if the boy sits n meters, the girls sits 3/2 n meters, and their sum has to be equal to 6 n + 3/2n = 6 find n, and you can define their distances from the pivot

OpenStudy (anonymous):

so, n +3/2n =6 n=6/3/2 n=4